multiply
(-1+2i)^2.
& if you can could you please show me how to plot the product on the complex plane?...if you can.

Answer by stanbon(48536)

multiply
(-1+2i)^2.
& if you can could you please show me how to plot the product on the complex plane?
-----------
(-1+2i)(-1+2i)
Use FOIL to get:
F: -1*-1=1
O: -1*2i=-2i
I: -1*21=-2i
L; 2i*2i = -4
----------------
Combine these to get:
-3-4i
-----------
The complex plane plot of this is the point (-3,-4)
=============
Cheers,
Stan H.